# From tests to effect sizes: Quantifying uncertainty and statistical variability in multilingual and multitask NLP evaluation benchmarks

### v1

@@ C.1 Variance components
$\text{ara}_{\text{AQMAR}}$ 0.61 1.61 0.61 1.87 1.60 2.70 2.05 2.87 0.32 0.72 0.45 1.05
$\text{bam}_{\text{MasakhaNER\_2.0}}$ 0.77 1.48 0.79 1.98 1.46 2.11 1.06 2.02 0.27 0.73 0.19 0.33
$\text{bar}_{\text{BarNER}}$ 1.28 2.08 1.85 2.02 1.57 2.09 2.10 2.25 0.52 0.94 0.18 0.34
$\text{bbj}_{\text{MasakhaNER\_2.0}}$ 0.75 1.73 1.59 1.79 0.89 1.94 0.49 0.83 0.68 0.93 0.23 0.45
$\text{cat}_{\text{AnCora}}$ 0.82 1.72 0.81 1.67 0.29 0.12 0.58 0.84 0.52 1.22 0.37 0.45
$\text{cmn}_{\text{UNER\_Chinese\_GSD}}$ 1.16 1.57 0.88 1.19 1.97 0.12 0.34 0.94 0.79 1.08 0.14 0.34
$\text{cmn}_{\text{UNER\_Chinese\_GSDSIMP}}$ 0.76 1.15 0.30 1.14 0.11 0.21 0.57 0.97 0.34 0.93 0.21 0.37
$\text{dan}_{\text{DaNE}}$ 1.42 1.18 0.88 1.23 0.06 0.13 0.41 0.84 0.58 0.96 1.41 2.43
$\text{deu}_{\text{GermEval}}$ 0.95 1.35 0.59 1.27 0.07 0.13 0.59 0.92 1.98 3.06 0.61 3.07
$\text{ell}_{\text{elNER}}$ 0.84 1.26 0.74 1.17 0.04 0.12 0.41 1.06 0.61 4.24 2.18 3.14
$\text{eng}_{\text{Tweebank}}$ 0.93 1.21 0.66 1.13 0.32 0.56 0.60 1.09 1.20 2.96 0.61 3.28
$\text{eng}_{\text{UNER\_English\_EWT}}$ 0.83 1.32 0.78 1.49 0.27 0.53 0.67 1.34 1.11 2.50 1.07 2.41
$\text{eus}_{\text{EIEC}}$ 1.54 2.81 0.92 1.44 0.41 0.74 0.56 1.16 2.40 2.66 1.86 2.45
$\text{ewe}_{\text{MasakhaNER\_2.0}}$ 2.63 3.29 0.71 1.93 0.38 0.69 0.27 1.12 0.93 3.08 0.26 0.56
$\text{fin}_{\text{TurkuNLP}}$ 2.10 3.68 0.83 1.69 0.24 0.57 0.47 1.08 0.29 0.52 0.22 0.51
$\text{fon}_{\text{MasakhaNER\_2.0}}$ 1.47 3.72 0.85 1.41 0.39 0.56 0.89 1.41 0.35 0.49 0.28 0.63
$\text{glg}_{\text{SLI\_Galician\_Corpora}}$ 1.29 2.80 0.89 1.47 0.65 0.88 1.30 1.43 0.40 0.66 0.25 0.69
$\text{hau}_{\text{MasakhaNER}}$ 2.02 3.27 0.42 0.59 1.40 0.95 0.77 1.67 0.29 0.54 0.24 0.55
$\text{hau}_{\text{MasakhaNER\_2.0}}$ 0.80 1.56 0.46 0.67 0.72 1.03 1.24 1.52 0.18 0.48 0.23 0.52
$\text{heb}_{\text{NEMO\_SPMRL}}$ 1.04 1.64 0.54 0.70 0.79 1.29 0.55 1.36 0.19 0.47 0.56 0.80
$\text{heb}_{\text{NEMO\_UD}}$ 1.23 1.55 0.43 0.79 0.60 0.93 1.21 1.46 0.72 1.29 0.77 0.86
$\text{hin}_{\text{HiNER}}$ 1.15 1.54 0.32 0.60 0.96 0.97 0.27 0.54 1.64 1.64 0.64 0.96
$\text{hrv}_{\text{hr500k}}$ 0.92 1.62 0.45 0.70 0.20 0.43 0.27 0.56 0.96 1.74 1.26 1.09
$\text{ibo}_{\text{MasakhaNER}}$ 0.90 1.76 0.76 1.27 0.91 0.52 0.40 1.03 1.67 1.68 0.66 0.87
$\text{ibo}_{\text{MasakhaNER\_2.0}}$ 0.24 0.72 1.13 1.34 0.60 0.63 0.36 0.72 0.72 1.26 0.84 0.93
$\text{kaz}_{\text{KazNERD}}$ 0.67 0.72 0.68 1.60 0.55 0.79 0.31 0.55 1.19 1.56 0.62 1.35
$\text{kin}_{\text{MasakhaNER}}$ 0.41 0.99 0.83 1.43 0.22 0.46 0.36 0.58 0.35 0.73 0.55 1.51
$\text{kin}_{\text{MasakhaNER\_2.0}}$ 0.40 0.81 0.56 1.37 0.82 0.55 0.29 0.58 0.42 0.80 0.91 1.74
$\text{lug}_{\text{MasakhaNER}}$ 0.46 0.72 0.85 1.23 0.26 0.61 0.41 0.61 0.54 1.16 0.75 1.80
$\text{lug}_{\text{MasakhaNER\_2.0}}$ 0.34 0.67 0.68 1.17 2.34 0.46 0.93 1.03 0.59 0.89 0.44 1.47
$\text{luo}_{\text{MasakhaNER}}$ 0.86 1.40 0.70 1.30 0.57 1.06 0.40 0.71 0.53 0.69 1.08 1.58
$\text{luo}_{\text{MasakhaNER\_2.0}}$ 0.61 1.56 0.92 1.38 0.39 0.50 0.19 0.61 0.44 0.65 0.83 2.52
$\text{mar}_{\text{L3Cube\_MahaNER}}$ 1.08 2.70 1.05 1.37 0.37 0.60 0.34 0.58 0.84 4.77 0.61 2.70
$\text{mos}_{\text{MasakhaNER\_2.0}}$ 0.46 1.42 0.74 1.29 0.61 0.42 0.39 1.14 1.14 1.07 1.19 2.67
$\text{nep}_{\text{EverestNER}}$ 0.78 1.44 0.47 1.29 0.73 1.31 0.54 1.13 0.91 1.07 1.22 2.85
$\text{nld}_{\text{CONLL02}}$ 0.56 1.61 0.75 1.02 0.88 1.57 1.28 1.65 1.71 1.07 1.11 2.48
$\text{nno}_{\text{NorNE}}$ 0.79 1.56 0.56 1.08 1.04 1.56 0.68 1.42 0.93 1.07 1.42 2.76
$\text{nob}_{\text{NorNE}}$ 0.59 1.61 0.81 1.18 2.06 1.68 0.46 1.15 1.57 1.07 0.28 0.83
$\text{nya}_{\text{MasakhaNER\_2.0}}$ 1.02 3.07 0.61 1.14 0.57 1.37 0.79 1.13 0.26 0.40 0.56 0.94
$\text{pcm}_{\text{MasakhaNER}}$ 0.47 1.52 0.75 0.99 1.21 1.47 0.32 0.95 0.26 0.44 0.51 1.03
$\text{pcm}_{\text{MasakhaNER\_2.0}}$ 0.65 1.56 0.38 1.06 0.30 0.57 0.80 1.02 0.39 0.55 0.57 1.05
$\text{por}_{\text{UNER Portuguese}}$ 0.59 1.62 0.23 1.08 0.48 0.67 1.57 1.52 0.49 0.68 0.44 0.99
$\text{qaf}_{\text{UNER Arabizi}}$ 1.22 2.00 0.43 1.02 0.32 0.71 0.97 1.31 0.33 0.40 1.10 1.06
$\text{ron}_{\text{RONEC}}$ 1.47 1.99 0.87 1.26 0.49 0.90 0.53 0.92 0.48 0.64 0.32 0.61
$\text{slk}_{\text{UNER\_Slovak\_SNK}}$ 1.50 2.36 0.49 1.22 0.25 0.64 0.68 0.99 0.33 0.54 0.32 0.65
$\text{slk}_{\text{WikiGoldSK}}$ 1.05 1.98 0.52 1.08 0.71 0.71 0.38 0.47 0.15 0.47 0.43 0.84
$\text{slv}_{\text{ssj500k}}$ 1.17 2.02 0.61 0.91 0.65 1.85 0.27 0.46 0.34 0.64 0.43 0.94
$\text{sna}_{\text{MasakhaNER\_2.0}}$ 0.65 1.71 0.48 0.64 1.68 2.02 0.47 0.60 0.26 0.55 0.38 0.62
$\text{spa}_{\text{AnCora}}$ 0.18 0.55 0.69 0.67 0.68 1.98 0.65 0.65 0.22 0.53 0.63 0.68
$\text{spa}_{\text{CONLL02}}$ 0.18 0.47 0.65 0.86 1.21 1.81 0.19 0.51 0.22 0.46 0.44 1.62
$\text{swa}_{\text{MasakhaNER}}$ 0.38 0.67 1.07 0.97 0.65 1.83 0.33 0.53 0.24 0.68 0.61 1.62
$\text{swa}_{\text{MasakhaNER\_2.0}}$ 0.30 0.52 0.46 0.65 1.29 1.97 0.37 0.97 0.43 0.68 0.50 1.66
$\text{swe}_{\text{UNER\_Swedish\_Talkbanken}}$ 0.22 0.53 0.61 0.66 0.31 0.56 0.86 1.15 0.28 0.70 0.58 1.48
$\text{tha}_{\text{ThaiNNER}}$ 0.29 0.50 0.57 1.42 0.22 0.57 0.67 1.25 0.19 0.70 0.64 1.71
$\text{tsn}_{\text{MasakhaNER\_2.0}}$ 0.29 0.67 1.29 1.86 0.48 0.66 0.96 1.51 0.50 0.68 1.11 1.61
$\text{twi}_{\text{MasakhaNER\_2.0}}$ 0.24 0.62 0.73 2.18 0.65 0.74 0.62 1.11 0.38 0.63 0.43 0.79
$\text{wol}_{\text{MasakhaNER}}$ 0.45 0.88 1.63 1.75 0.41 0.56 1.70 1.33 0.45 0.91 0.86 0.79
$\text{wol}_{\text{MasakhaNER\_2.0}}$ 0.37 0.72 0.73 1.43 0.29 0.64 0.37 0.67 0.48 1.09 0.75 1.02
$\text{xho}_{\text{MasakhaNER\_2.0}}$ 0.27 0.64 0.68 1.68 0.77 2.16 0.50 0.69 0.44 1.12 0.57 0.88
$\text{yor}_{\text{MasakhaNER}}$ 0.13 0.66 1.12 2.14 0.61 2.79 0.65 0.88 0.68 1.22 0.68 0.81
$\text{yor}_{\text{MasakhaNER\_2.0}}$ 1.63 1.75 1.60 2.15 1.25 2.24 0.75 0.81 0.49 0.91 0.52 0.82
@@
C.2 Aggregate inferences
Table 13 shows uncertainty quantified estimates of aggregate performance for the NER experiments.
Model Statistic F1 (true) Estimate SE 95% CI (percentile) 95% CI ($\hat{\theta}\pm$ 2SE) CI ($\hat{\theta}\pm$ IQW)
Glot500 (c.) $\hat{\mu}_{\text{am}}$ 85.62 85.61 0.20 85.21 85.99 85.21 86.01 85.22 86.00
$\hat{\mu}_{\text{gm}}$ 85.62 85.35 0.22 84.90 85.77 84.91 85.79 84.92 85.79
$\mu_{\text{md}}$ 85.62 87.53 0.38 86.73 88.23 86.77 88.29 86.78 88.28
Glot500 (i.) $\hat{\mu}_{\text{am}}$ 83.71 83.71 0.41 82.82 84.40 82.89 84.52 82.91 84.50
$\hat{\mu}_{\text{gm}}$ 83.71 83.17 0.55 81.90 84.04 82.07 84.27 82.10 84.24
$\mu_{\text{md}}$ 83.71 86.46 0.42 85.63 87.28 85.62 87.30 85.64 87.29
mBERT (c.) $\hat{\mu}_{\text{am}}$ 72.60 72.61 0.24 72.13 73.07 72.12 73.09 72.13 73.08
$\hat{\mu}_{\text{gm}}$ 72.60 70.57 0.30 69.97 71.16 69.96 71.18 69.98 71.17
$\mu_{\text{md}}$ 72.60 76.34 0.49 75.36 77.24 75.37 77.31 75.40 77.28
mBERT (i.) $\hat{\mu}_{\text{am}}$ 77.31 77.31 0.40 76.40 77.92 76.52 78.11 76.56 78.07
$\hat{\mu}_{\text{gm}}$ 77.31 76.51 0.58 75.10 77.27 75.35 77.67 75.42 77.60
$\mu_{\text{md}}$ 77.31 79.61 0.48 78.70 80.53 78.65 80.57 78.69 80.52
XLM-R (c.) $\hat{\mu}_{\text{am}}$ 84.91 84.88 0.20 84.47 85.28 84.47 85.29 84.48 85.29
$\hat{\mu}_{\text{gm}}$ 84.91 84.62 0.23 84.17 85.06 84.16 85.07 84.17 85.06
$\mu_{\text{md}}$ 84.91 86.71 0.39 85.95 87.47 85.94 87.49 85.95 87.47
XLM-R (i.) $\hat{\mu}_{\text{am}}$ 83.88 83.87 0.49 82.41 84.43 82.88 84.85 82.85 84.88
$\hat{\mu}_{\text{gm}}$ 83.88 83.28 1.19 79.59 84.14 80.89 85.66 81.01 85.55
$\mu_{\text{md}}$ 83.88 86.17 0.52 85.09 87.14 85.13 87.22 85.15 87.20
Glot500 (c.) $\hat{\mu}_{\text{am}}$ 85.62 85.85 0.58 84.72 87.01 84.69 87.01 84.70 86.99
$\hat{\mu}_{\text{gm}}$ 85.62 85.60 0.62 84.39 86.85 84.37 86.84 84.37 86.83
$\mu_{\text{md}}$ 85.62 87.67 0.66 86.16 88.75 86.35 88.99 86.37 88.96
Glot500 (i.) $\hat{\mu}_{\text{am}}$ 83.71 83.81 0.84 82.12 85.39 82.14 85.49 82.18 85.44
$\hat{\mu}_{\text{gm}}$ 83.71 83.27 1.01 81.10 85.07 81.25 85.29 81.28 85.25
$\mu_{\text{md}}$ 83.71 86.55 0.74 84.96 87.89 85.06 88.04 85.08 88.01
mBERT (c.) $\hat{\mu}_{\text{am}}$ 72.60 73.58 1.30 71.07 76.14 70.97 76.18 71.04 76.12
$\hat{\mu}_{\text{gm}}$ 72.60 71.67 1.71 68.47 75.18 68.25 75.10 68.32 75.03
$\mu_{\text{md}}$ 72.60 76.99 0.95 75.12 78.95 75.10 78.89 75.08 78.91
mBERT (i.) $\hat{\mu}_{\text{am}}$ 77.31 77.24 0.93 75.32 79.05 75.38 79.10 75.37 79.11
$\hat{\mu}_{\text{gm}}$ 77.31 76.45 1.10 74.07 78.52 74.25 78.65 74.23 78.67
$\mu_{\text{md}}$ 77.31 79.57 0.80 78.03 81.16 77.96 81.17 78.00 81.13
XLM-R (c.) $\hat{\mu}_{\text{am}}$ 84.91 85.05 0.57 83.95 86.16 83.91 86.18 83.94 86.15
$\hat{\mu}_{\text{gm}}$ 84.91 84.79 0.60 83.63 85.99 83.58 86.00 83.61 85.97
$\mu_{\text{md}}$ 84.91 86.79 0.68 85.40 88.06 85.43 88.16 85.46 88.12
XLM-R (i.) $\hat{\mu}_{\text{am}}$ 83.88 83.84 0.82 81.78 85.24 82.20 85.49 82.11 85.57
$\hat{\mu}_{\text{gm}}$ 83.88 83.27 1.52 78.07 84.98 80.23 86.32 79.82 86.73
$\mu_{\text{md}}$ 83.88 86.03 0.92 84.12 87.60 84.19 87.87 84.29 87.77
Table 13: Overall model performance with languages fixed (top) and randomized (bottom).
C.3 Ranks

### v2

@@ C.1 Variance components
$\text{ara}_{\text{AQMAR}}$ 0.44 2.63 0.71 2.21 1.81 4.37 0.62 3.04 0.83 2.97 1.54 2.50
$\text{bam}_{\text{MasakhaNER 2.0}}$ 0.64 1.34 2.22 1.41 1.10 1.85 0.74 1.62 0.90 1.48 0.68 1.75
$\text{bar}_{\text{BarNER}}$ 2.52 7.97 7.43 10.91 4.57 13.64 2.34 13.96 1.89 7.94 4.27 10.80
$\text{bbj}_{\text{MasakhaNER 2.0}}$ 0.66 2.46 1.16 2.73 1.54 2.43 1.45 2.38 1.03 2.66 0.86 3.13
$\text{cat}_{\text{AnCora}}$ 0.06 0.53 0.50 0.53 0.18 0.98 0.16 0.66 0.24 0.52 0.14 0.45
$\text{cmn}_{\text{UNER Chinese GSDSIMP}}$ 0.70 2.45 0.41 2.60 1.13 9.52 0.27 2.33 0.50 2.47 0.40 2.66
$\text{cmn}_{\text{UNER Chinese GSD}}$ 0.73 1.99 72.48 2.45 1.25 7.35 0.22 2.04 0.63 2.11 0.28 2.62
$\text{dan}_{\text{DaNE}}$ 1.83 4.02 2.67 4.01 2.53 5.61 1.12 3.96 1.49 4.10 0.49 2.96
$\text{deu}_{\text{GermEval}}$ 0.04 0.30 0.04 0.23 0.16 0.46 0.10 0.27 0.06 0.28 0.09 0.26
$\text{ell}_{\text{elNER}}$ 0.10 0.46 0.06 0.39 0.22 0.79 0.15 0.52 0.08 0.41 0.02 0.43
$\text{eng}_{\text{Tweebank}}$ 2.66 3.11 61.11 3.53 0.68 3.97 3.46 4.13 2.66 3.23 0.74 2.81
$\text{eng}_{\text{UNER English EWT}}$ 0.87 1.42 0.10 1.31 0.81 1.52 0.40 1.62 0.54 1.39 0.51 1.29
$\text{eus}_{\text{EIEC}}$ 0.66 2.24 1.00 2.09 0.57 3.75 0.75 2.88 0.78 2.01 0.83 2.19
$\text{ewe}_{\text{MasakhaNER 2.0}}$ 0.19 0.35 0.24 0.45 0.33 0.49 0.19 0.63 0.10 0.37 0.23 0.50
$\text{fin}_{\text{TurkuNLP}}$ 0.68 1.62 1.35 1.82 0.57 2.58 0.73 2.08 0.39 1.89 0.80 1.53
$\text{fon}_{\text{MasakhaNER 2.0}}$ 0.54 1.39 0.54 1.70 0.88 1.93 1.21 1.90 0.62 1.68 0.24 1.68
$\text{glg}_{\text{SLI Galician Corpora}}$ 0.58 1.04 0.34 1.18 0.69 1.40 0.43 1.32 0.64 1.00 0.19 1.13
$\text{hau}_{\text{MasakhaNER 2.0}}$ 0.26 0.42 0.52 0.45 0.46 0.74 1.34 0.95 0.23 0.43 0.42 0.44
$\text{hau}_{\text{MasakhaNER}}$ 0.07 1.17 0.21 1.04 0.93 1.61 0.31 1.50 0.29 1.17 0.40 0.84
$\text{heb}_{\text{NEMO SPMRL}}$ 0.33 2.03 1.79 3.48 0.68 4.80 3.01 3.10 0.60 2.05 0.51 2.84
$\text{heb}_{\text{NEMO UD}}$ 1.45 4.60 2.55 4.69 2.86 7.36 2.30 4.51 2.73 4.42 0.90 3.78
$\text{hin}_{\text{HiNER}}$ 0.09 0.02 4.34 0.01 0.01 0.04 0.00 0.02 0.01 0.02 0.00 0.01
$\text{hrv}_{\text{hr500k}}$ 0.11 0.32 0.08 0.28 0.17 0.55 0.15 0.47 0.06 0.32 0.17 0.32
$\text{ibo}_{\text{MasakhaNER 2.0}}$ 0.04 0.19 0.92 0.27 0.39 0.40 0.35 0.62 0.05 0.22 0.81 0.31
$\text{ibo}_{\text{MasakhaNER}}$ 0.45 0.78 2.13 0.91 0.63 1.06 0.66 1.68 0.34 0.87 0.99 0.96
$\text{kaz}_{\text{KazNERD}}$ 0.07 0.38 6.27 0.21 0.36 1.14 0.18 0.25 0.15 0.36 790.74 0.18
$\text{kin}_{\text{MasakhaNER 2.0}}$ 0.09 0.33 0.25 0.45 0.11 0.52 0.27 0.81 0.07 0.42 0.55 0.51
$\text{kin}_{\text{MasakhaNER}}$ 0.58 1.73 0.90 2.49 1.23 2.46 4.51 2.86 0.37 1.88 1.71 2.20
$\text{lug}_{\text{MasakhaNER 2.0}}$ 0.10 0.32 0.06 0.33 0.25 0.44 0.44 0.56 0.20 0.32 0.10 0.42
$\text{lug}_{\text{MasakhaNER}}$ 0.46 3.47 3.27 4.14 0.51 3.97 1.58 3.32 0.48 3.38 1.82 3.92
$\text{luo}_{\text{MasakhaNER 2.0}}$ 0.27 0.70 0.39 0.72 0.11 0.90 0.34 0.94 0.20 0.72 0.40 0.85
$\text{luo}_{\text{MasakhaNER}}$ 0.68 4.71 187.53 7.88 1.72 5.05 4.92 8.33 1.23 4.10 4.90 5.14
$\text{mar}_{\text{L3Cube MahaNER}}$ 0.19 1.13 0.38 1.19 0.51 1.82 0.33 1.35 0.08 1.28 0.25 1.18
$\text{mos}_{\text{MasakhaNER 2.0}}$ 0.83 2.00 1.91 2.06 0.62 2.83 1.77 2.35 0.33 1.86 1.72 2.16
$\text{nep}_{\text{EverestNER}}$ 0.07 0.29 0.09 0.32 0.19 1.07 0.15 0.53 0.09 0.31 0.14 0.34
$\text{nld}_{\text{CONLL02}}$ 0.08 0.33 0.19 0.38 0.95 1.07 0.16 0.51 0.04 0.37 0.14 0.35
$\text{nno}_{\text{NorNE}}$ 0.15 1.31 0.30 1.28 1.91 2.74 0.52 2.03 0.22 1.33 0.70 1.29
$\text{nob}_{\text{NorNE}}$ 0.14 0.92 0.77 1.06 2.95 2.33 1.01 1.73 0.30 0.86 0.55 0.99
$\text{nya}_{\text{MasakhaNER 2.0}}$ 0.16 0.22 0.07 0.21 0.25 0.36 0.47 0.42 0.05 0.26 0.12 0.29
$\text{pcm}_{\text{MasakhaNER 2.0}}$ 0.15 0.45 0.30 0.49 0.45 0.79 0.65 0.67 0.11 0.52 0.09 0.54
$\text{pcm}_{\text{MasakhaNER}}$ 0.14 0.96 0.75 1.33 0.50 1.59 1.03 2.30 0.44 1.25 3.29 1.77
$\text{por}_{\text{UNER Portuguese}}$ 0.29 0.90 0.50 0.88 0.28 1.50 0.77 1.17 0.12 0.88 0.40 0.94
$\text{qaf}_{\text{UNER Arabizi}}$ 4.37 9.44 27.94 18.19 1.55 8.86 1.49 6.32 6.06 7.12 1.01 9.57
$\text{ron}_{\text{RONEC}}$ 0.09 0.27 0.13 0.25 0.17 0.44 0.09 0.29 0.04 0.23 0.04 0.22
$\text{slk}_{\text{UNER Slovak SNK}}$ 0.56 1.67 2.92 2.71 0.94 3.07 3.14 2.86 0.50 1.61 1.57 2.46
$\text{slk}_{\text{WikiGoldSK}}$ 0.14 0.54 0.18 0.64 0.34 1.35 0.37 0.80 0.32 0.49 0.22 0.42
$\text{slv}_{\text{ssj500k}}$ 0.82 22.96 1.37 28.19 0.99 28.37 3.54 27.08 0.86 24.95 2.84 27.24
$\text{sna}_{\text{MasakhaNER 2.0}}$ 0.08 0.16 0.06 0.19 0.17 0.30 0.22 0.46 0.10 0.16 0.22 0.41
$\text{spa}_{\text{AnCora}}$ 0.11 0.29 0.03 0.22 0.12 0.42 0.07 0.31 0.05 0.28 0.05 0.22
$\text{spa}_{\text{CONLL02}}$ 0.06 0.47 0.21 0.46 0.09 0.50 0.04 0.49 0.28 0.46 0.16 0.40
$\text{swa}_{\text{MasakhaNER 2.0}}$ 0.04 0.11 0.04 0.12 0.06 0.20 0.14 0.20 0.02 0.12 0.05 0.14
$\text{swa}_{\text{MasakhaNER}}$ 0.20 0.83 0.27 1.20 0.23 1.26 0.54 1.51 0.26 0.84 0.26 1.12
$\text{swe}_{\text{UNER Swedish Talkbanken}}$ 2.10 5.94 102.29 9.49 5.23 9.95 414.24 10.89 1.40 5.85 3.88 6.05
$\text{tha}_{\text{ThaiNNER}}$ 0.07 0.31 0.06 0.27 0.08 0.40 0.06 0.48 0.06 0.31 0.06 0.27
$\text{tsn}_{\text{MasakhaNER 2.0}}$ 0.35 0.64 0.65 0.76 0.50 0.93 1.74 1.21 0.48 0.76 0.78 0.88
$\text{twi}_{\text{MasakhaNER 2.0}}$ 0.40 1.84 0.38 2.30 0.93 3.04 0.57 3.29 0.19 2.18 1.39 2.52
$\text{wol}_{\text{MasakhaNER 2.0}}$ 0.09 0.69 0.39 0.90 0.31 1.06 0.40 1.11 0.20 0.99 1.39 1.13
$\text{wol}_{\text{MasakhaNER}}$ 0.73 6.40 17.24 7.39 1.60 7.20 1.65 8.22 1.31 6.19 2.26 7.67
$\text{xho}_{\text{MasakhaNER 2.0}}$ 0.12 0.37 0.11 0.43 0.20 0.71 0.20 0.90 0.17 0.39 0.44 0.46
$\text{yor}_{\text{MasakhaNER 2.0}}$ 0.20 0.63 0.83 0.63 0.65 1.05 0.35 0.78 0.51 0.66 0.32 0.68
$\text{yor}_{\text{MasakhaNER}}$ 0.21 2.65 8.92 2.65 0.27 2.80 0.38 2.21 0.48 2.94 1.29 2.61
@@
C.2 Ranks